The record for the Boston marathons wheelchair division is 1 hour,18 minutes and 27 seconds. The Boston marathon is 26.2 miles long. What was the average speed of the record winner of the wheelchair division?
Answer by stanbon(75887)

The record for the Boston marathons wheelchair division is 1 hour,18 minutes and 27 seconds. The Boston marathon is 26.2 miles long. What was the average speed of the record
-----------------
Convert to hours: 1 hr + (18/60)hr + (27/3600)hr = 1.3075 hrs
------
rate = distance/time = 26.2/1.3075 mph = 20.04 mph
============================
Cheers,
Stan H.
============================